A rectangular coil of single turn, having area A, rotates in a uniform magnetic field B an angular velocity w about an axis is perpendicular to the field .If initially the plane of coil is perpendicular to the field,then the average induced e.m.f. when it has rotated through 90 degree is???a. wBA/pie b. wBA/2 pie c. wBA/4 pie d. 2wBA/ pie


Grade:upto college level

1 Answers

ROSHAN MUJEEB
3 years ago
initially flux,ϕ=BAcos0=BA
After rotating through an angle90o.
Flux through the coil is zero.
So,Δϕ=BA

Angular speed=w, so, time period=ω/2π​=T

4T​is time taken to rotate90o.

So,ΔtΔϕ​=T/4BA​=π2BAω​